Fermi National Laboratory


Accelerator Update for January 24, 2005 - January 26, 2005

Monday January 24
The day shift began with cryo system technicians working on the Tevatron (TeV) E2 wet engine, and with MiniBooNE and the Meson experiment MCenter taking beam.

The Neutron Therapy Facility began taking bursts of beam at 2:02 PM.

Main Injector (MI) experts began a study period at 2:15 PM.

Operators started the TeV check out at 2:24 PM.

At 3 PM, a Booster kicker (MKS02) tripped off. An expert investigated.

Operations put the TeV into a dry squeeze at 4:30 PM.

At 5:19 PM, the TeV sector E2 cold compressor tripped off, which was followed by a TeV quench. Cryo system techs at CHL investigated and then reported that they would replace the E2 cold compressor.

Operators prepared the Collider Detector at Fermilab (CDF) and D-Zero (D0) collision halls for access at 5:42 PM.

Operations reported at 10:44 PM that they had checked Booster coalescing and then began tuning.

Tuesday January 25
The midnight shift began with the TeV in a dry squeeze, with Booster tuning continuing, and with MiniBooNE taking beam.

Operations put the TeV in shot setup at 12:27 AM and disabled the Switchyard (SY) beam.

At 3:18 AM, Operations reported having problems with loading the TeV with protons. Both a TeV and a Booster expert were helping solve the problems.

Operations reported that protons had been loaded into the TeV by 4:42 AM.

Operations established store 3945 at 5:29 AM with an initial luminosity of 99.5E30. Pbar resumed stacking.

Operations reestablished beam to MCenter at 6 AM.

At 9:17 AM, Operations reported having a problem getting the antiprotons from the Debuncher into the Accumulator due to a kicker (A:IKIK) not firing. An expert investigated and began repairs. The Run Coordinator decided to use this time to stash antiprotons into the Recycler.

At 11:38 AM, two Booster RF stations (BRF1 & 2) tripped off for the first of four times in two hours. The trips were due to vacuum high pressure. Each time, the stations reset okay.

The shots to the Recycler were completed by 12:55 PM.

At 1:40 PM, BRF1 & 2 tripped off again. They reset okay. An expert investigated and found a bad cable.

Operations reported at 2:39 PM that the Accumulator kicker was repaired.

At 2:54 PM, Operations reported that an Accumulator RF station (ARF1) wouldn't turn on. An expert investigated.

Stacking resumed at 3:36 PM, but then stopped two minutes later with more ARF1 problems.

At 3:58 PM, BRF1 tripped off and wouldn't reset. An expert got it on ten minutes later.

Stacking resumed at 4:07 PM.

At 7:43 PM, a Linac RF station (KRF7) had a regulation problem due to a temperature problem. An expert came in and worked on the station's LCW regulation loops.

At 11:21 PM, Pbar began having more ARF1 problems.

Wednesday January 26
The midnight shift began with operations monitoring store 3945, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

Around 2:30 AM, a cryo system compressor tripped off. Operators talked with the CHL Crew Chief; he said to leave the compressor off.

At 3:47 AM, Operations reported having more problems with ARF1. Operators made adjustments and resets throughout the rest of the shift to keep the station running.

Operations disabled the SY beam at 7:01 AM.

Operations began turning off MI at 7:05 AM for the planned MI-60 VCB cubical removal.

The Plan for Wednesday
The plan for today is to wait until the MI-60 cubical work is finished, which should be around noon. Operations will try to keep the present store in the machine until this evening and then establish a new store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/26/2005   email Fermilab